Grade 11Electrostatics

a pendulem bob of mass 30mg carrying a charge of 2*10^-8 is at rest in auniform horiontal electric field of 20000N/c.the tension in the thread of the pendulem is

Prasanna Parthiban
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
ere we have given that mass is 30mg 
What's more, the charge i.e. q = 2*10^-8 and E = 20000N/c 
Give T a chance to be the pressure and Ѳ be the edge with the vertical. 
Presently in balance 
Tsin Ѳ = qE and Tcos Ѳ=mg 
At that point tan Ѳ = qE/mg 
=(2*10^-8 x20000)/(30x 10^-6 x9.8) 
=(4x 10^-4)/(294x10^-6) 
= 0.0136 x10^2 
Ѳ = 53.7o 
What's more, we additionally realize that Tsin Ѳ = qE 
T = qE/sinѳ 
= (2*10^-8 x20000)/sin53.7 
=(4x10^-4)/0.806 
= 4.96 x 10^-4 N